OpenStudy (anonymous):

What is the vertex of the graph of y = –1(x – 2)2 + 3?

OpenStudy (anonymous):

maybe.. not completely sure.. i suck at math

OpenStudy (anonymous):

do you know what a vertex is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Vertiex: -The highest point; the top or apex. (google dictionary) do looking at the graph i sent you, what is the highest point?

OpenStudy (anonymous):

vertex*

OpenStudy (anonymous):

2,3 ?

OpenStudy (anonymous):

yes! good job:)

OpenStudy (anonymous):

AHHH thankyou!!!!!!!!!(((((::

OpenStudy (anonymous):

can you help me with the rest of my awful test?

OpenStudy (anonymous):

hahah no problem:) glad i could help

OpenStudy (anonymous):

um sure.

OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 – 10x + 21?

OpenStudy (anonymous):

http://mathway.com/graph_image.aspx?p=y=x2-10x+21?p=520?p=450?p=False?p=False?p=False?p=False?p=True?p=False?p=True?p=True?p=?p=?p=?p=?p=4 theres the graph. were there supposed to be things after x- and y- ?

OpenStudy (anonymous):

oh just kidding.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

forget the question i asked. um so you see the graph? do you know what x and y intercepts are?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

y = –1(x – 2)2 + 3 y=(-x+2)2 + 3 y= -x^2 + 4 + 3 y= -x^2 + 7 y= x^2 - 7 so thats the y intercept. use this ==> y = –1(x – 2)2 + 3 and solve for x to get the x intercept

OpenStudy (anonymous):

y=(1x)5 ?

OpenStudy (anonymous):

but you're solving for x:) so it should look like this x= ______

OpenStudy (anonymous):

idk how id do that...

OpenStudy (anonymous):

y = –1(x – 2)2 + 3 y=(-x+2)2 + 3 y= -x^2 + 4 + 3 y= -x^2 + 7 y= x^2 - 7 good so far? its what i did earlier. y= x^2 - 7 7+y=x^2 x= \[\sqrt{7+y}\]

OpenStudy (anonymous):

i might be wrong... eek. i think i am.

OpenStudy (anonymous):

i'll just guess, i guess..and I can ask you my next question?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Which statement describes the translation of y = – (x – 2)2 – 2 from standard position? these are the answers to choose from Moved up and to the left. Moved up and to the right. Moved down and to the left. Moved down and to the right.

OpenStudy (anonymous):

Heyyy i got the one i though i was wrong on. x= http://mathway.com/math_image.aspx?p=(SMB02RSMB033SMB02rSMB03+2,0),(-SMB02RSMB033SMB02rSMB03+2,0)?p=167?p=22 y=(0,-1) and for the last question it's : to the right.

OpenStudy (anonymous):

moved up or down?

OpenStudy (anonymous):

not sure. i'm guessing down

OpenStudy (anonymous):

What is the range of the graph of y = 5(x – 2)2 + 7? Answer y 2 y 7 y 7 y 2

OpenStudy (anonymous):

*** \[y \ge 2\] \[y \le 7\] \[y \ge 7\] \[y \le 2\]

OpenStudy (anonymous):

What is the equation of the graph below in vertex form, y=a(x – h)2 + k? Using complete sentences, explain how you developed your equation

OpenStudy (anonymous):

i actually got y≥7/5 so. i dont know.. i'd go for y≥2-- sorry about that. um, i feel like i'm doing your test for you.. which is cheating. so why dont you give it a try.

OpenStudy (anonymous):

ok..